# Constant Rate Scheduling: Constant-Rate Distributional Change for Efficient Training and Sampling in Diffusion Models

### v3

@@ Appendix G Generated Samples
Table 24: Evaluation results on FFHQ 64 $\times$ 64 using the pretrained model provided in EDM. Bold values indicate best sampling schedules for each combination of sampler and NFE. FID scores on AYS is taken from their original paper [16].
Sampling FID $\downarrow$
Sampler schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=50$
DPM-Solver++(2M) EDM 5.57 3.02 2.58
AYS 5.43 3.29 2.62
CRS-$v_{x}+v_{\mathrm{FID}}$ 4.70 2.87 2.52
iPNDM EDM 4.60 2.62 2.49
CRS-$v_{x}+v_{\mathrm{FID}}$ 3.87 2.53 2.44
Table 25: Evaluation results for CIFAR10 32 $\times$ 32 on class-conditional image generation using the pretrained model provided in EDM. Bold values indicate best sampling schedules for each combination of sampler and NFE.
Sampling FID $\downarrow$
Sampler schedule $\mathrm{NFE}=10$ $\mathrm{NFE}=20$ $\mathrm{NFE}=50$
DPM-Solver++(2M) EDM 4.12 2.20 1.94
CRS-$v_{x}+v_{\mathrm{FID}}$ 3.03 2.07 1.89
iPNDM EDM 2.51 1.93 1.89
CRS-$v_{x}+v_{\mathrm{FID}}$ 2.33 1.92 1.89
Figure 14: Samples of LSUN Horse 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=250$ ($\mathrm{FID}=2.03$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. SDE-DPM-Solver++(2M) was used as sampler.
Figure 15: Samples of LSUN Horse 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=50$ ($\mathrm{FID}=2.04$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. PNDM was used as sampler.
Figure 16: Samples of LSUN Bedroom 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=10$ ($\mathrm{FID}=3.30$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. UniPC was used as sampler.
Figure 17: Samples of LSUN Bedroom 256 $\times$ 256 generated using pixel-space diffusion model at $\mathrm{NFE}=5$ ($\mathrm{FID}=14.02$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. DPM-Solver++(2M) was used as sampler.
Figure 18: Samples of LSUN Church 256 $\times$ 256 generated using latent-space diffusion model at $\mathrm{NFE}=20$ ($\mathrm{FID}=3.89$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively. DPM-Solver++(2M) was used as sampler.
Figure 19: Samples of LSUN Bedroom 256 $\times$ 256 generated using latent-space diffusion model at $\mathrm{NFE}=20$ ($\mathrm{FID}=2.71$). We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively. PNDM was used as sampler.
Figure 20: Samples of ImageNet 256 $\times$ 256 generated using latent-space diffusion model at $\mathrm{NFE}=20$ with class condition ($\mathrm{FID}=10.83$). Six classes were randomly selected, and each row corresponds to one class. We optimized training and sampling schedules using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively. SDE-DPM-Solver++(2M) was used as sampler.

### v4

@@ Appendix G Generated Samples
Figure 14: Generated samples of LSUN Horse $\times256$ using a pixel-space diffusion model at $\mathrm{NFE}=250$ ($\mathrm{FID}=2.03$). The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. SDE-DPM-Solver++(2M) was used as the sampler.
Figure 15: Generated samples of LSUN Horse $\times256$ using a pixel-space diffusion model at $\mathrm{NFE}=50$ ($\mathrm{FID}=2.04$). The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. PNDM was used as the sampler.
Figure 16: Generated samples of LSUN Bedroom $\times256$ using a pixel-space diffusion model at $\mathrm{NFE}=10$ ($\mathrm{FID}=3.30$). The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. UniPC was used as the sampler.
Figure 17: Generated samples of LSUN Bedroom $\times256$ using a pixel-space diffusion model at $\mathrm{NFE}=5$ ($\mathrm{FID}=14.02$). The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{x}+v_{\mathrm{FID}}$, respectively. DPM-Solver++(2M) was used as the sampler.
Figure 18: Generated samples of LSUN Church $\times256$ using a latent-space diffusion model at $\mathrm{NFE}=20$ ($\mathrm{FID}=3.89$). The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively. DPM-Solver++(2M) was used as the sampler.
Figure 19: Generated samples of LSUN Bedroom $\times256$ using a latent-space diffusion model at $\mathrm{NFE}=20$ ($\mathrm{FID}=2.71$). The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively. PNDM was used as the sampler.
Figure 20: Generated samples of ImageNet $\times256$ using a latent-space diffusion model at $\mathrm{NFE}=20$ with class condition ($\mathrm{FID}=10.83$). Six classes were randomly selected, and each row corresponds to one class. The training and sampling schedules were optimized using CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, respectively. SDE-DPM-Solver++(2M) was used as the sampler.